OpenStudy (anonymous):

Form a sequence that has two arithmetic means between -18 and -144

OpenStudy (anonymous):

let a1 and a2 be two A.M's between -18 and -144 then -18,a1,a2,-144 are in A.P let d be c.d an=a+(n-1)d an=-144 a=-18 n=4 -144=-18+(4-1)d -144+18=3d find d then a1,a2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw